OpenStudy (anonymous):

The area of a rectangle is given by A = x2 + 4x – 5. Find the value of x if the total area of the rectangle is equal to 7 square units. a) x = 7 b) x = 1 c) x = 5 d) x = 2

OpenStudy (anonymous):

7=x^2+4x-5 0=x^2+4x-12 after this just factor and use the positive value